# Dynamical Formation of Apparent Horizons due to Boundary Effect in Vacuum Einstein Gravity

### v2

@@ 6.1. Energy Estimates for the Weyl/Riemann Curvature components
(176) $\begin{split}&\frac{1}{a^{\frac{1}{2}}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\alpha}\rVert_{L^{2}_{(sc)}(H_{u}^{(0,\underline{u})})}+\frac{1}{a^{\frac{1}{2}}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\beta}\rVert_{L^{2}_{(sc)}(\underline{H}_{\underline{u}}^{(u_{\infty},u)})}\\ \lesssim&\frac{1}{a^{\frac{1}{2}}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\alpha}\rVert_{L^{2}_{(sc)}(H_{u_{\infty}}^{(0,\underline{u})})}+\frac{1}{a^{\frac{1}{2}}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\beta}\rVert_{L^{2}_{(sc)}(\underline{H}_{0}^{(u_{\infty},u)})}\\ &+\frac{1}{a^{\frac{1}{2}}}\int_{0}^{\underline{u}}\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert}\lVert{a^{\frac{i}{2}}P_{i}\cdot(a^{\frac{1}{2}}\nabla)^{i}\alpha}\rVert_{L^{1}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}u^{\prime}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\\ &+\frac{1}{a^{\frac{1}{2}}}\int_{0}^{\underline{u}}\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert}\lVert{a^{\frac{i}{2}}Q_{i}\cdot(a^{\frac{1}{2}}\nabla)^{i}\beta}\rVert_{L^{1}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}u^{\prime}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\\ \lesssim&\frac{1}{a^{\frac{1}{2}}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\alpha}\rVert_{L^{2}_{(sc)}(H_{u_{\infty}}^{(0,\underline{u})})}+\frac{1}{a^{\frac{1}{2}}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\beta}\rVert_{L^{2}_{(sc)}(\underline{H}_{0}^{(u_{\infty},u)})}+\frac{1}{a^{\frac{1}{3}}}.\end{split}$
@@
Let $(\Psi_{1},\Psi_{2})\in\begin{Bmatrix}a^{\frac{1}{2}}(\beta,(\rho,\sigma)),a^{\frac{1}{2}}((\rho,\sigma),\underline{\beta}),(\underline{\beta},\underline{\alpha})\end{Bmatrix}$. For $0\leq i\leq N+4$, there holds:
(177) $\lVert{(a^{\frac{1}{2}}\nabla)^{i}\Psi_{1}}\rVert_{L^{2}_{(sc)}(H_{u}^{(0,\underline{u})})}+\lVert{(a^{\frac{1}{2}}\nabla)^{i}\Psi_{2}}\rVert_{L^{2}_{(sc)}(\underline{H}_{\underline{u}}^{(u_{\infty},u)})}$
(178) $\lesssim$ $\lVert{(a^{\frac{1}{2}}\nabla)^{i}\Psi_{1}}\rVert_{L^{2}_{(sc)}(H_{u_{\infty}}^{(0,\underline{u})})}+\lVert{(a^{\frac{1}{2}}\nabla)^{i}\Psi_{2}}\rVert_{L^{2}_{(sc)}(\underline{H}_{0}^{(u_{\infty},u)})}+1.$
@@
$\nabla_{3}\Psi_{1}+\left(\frac{1}{2}+s_{2}(\Psi_{1})\right)\mbox{tr}\underline{\chi}\Psi_{1}-\mathcal{D}\Psi_{2}=$ $(\psi,\hat{\chi})\Psi,$
(179) $\begin{split}\nabla_{4}\Psi_{2}-\prescript{*}{}{\mathcal{D}}\Psi_{1}=(\psi,\underline{\hat{\chi}})(\Psi_{u},\alpha).\end{split}$
@@
(180) $\begin{split}&\nabla_{3}\nabla^{i}\Psi_{1}+\left(\frac{i+1}{2}+s_{2}(\Psi_{1})\right)\mbox{tr}\underline{\chi}\nabla^{i}\Psi_{1}-\mathcal{D}\nabla^{i}\Psi_{2}\\ =&\sum_{i_{1}+i_{2}+i_{3}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}+1}\Psi_{2}+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\psi_{g},\hat{\chi})\nabla^{i_{4}}\Psi\\ &+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\underline{\hat{\chi}},\widetilde{\mbox{tr}\underline{\chi}})\nabla^{i_{4}}\Psi_{1}+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}\mbox{tr}\underline{\chi}\nabla^{i_{4}}\Psi_{1}\\ &+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-2}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}(\underline{\hat{\chi}},\mbox{tr}\underline{\chi})\nabla^{i_{4}}\Psi_{1}:=P_{i}.\end{split}$
@@
(181) $\begin{split}&\nabla_{4}\nabla^{i}\Psi_{2}-\prescript{*}{}{\mathcal{D}}\nabla^{i}\Psi_{1}\\ =&\sum_{i_{1}+i_{2}+i_{3}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}+1}\Psi_{1}+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\psi_{g},\underline{\hat{\chi}})\nabla^{i_{4}}(\Psi,\alpha)\\ &+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}\Psi_{2}\\ &+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-2}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}\Psi_{2}:=Q_{i}.\end{split}$
@@
(189) $\begin{split}&\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\left(\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-2}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}(\underline{\hat{\chi}},\mbox{tr}\underline{\chi})\nabla^{i_{4}}\Psi_{1}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{\frac{1}{2}}\hskip 1.42262pt\text{d}u^{\prime}\\ \lesssim&\frac{a}{\lvert u\rvert}\Gamma[\mbox{tr}\underline{\chi}]\Gamma[\alpha^{F}]^{2}\lesssim\frac{a}{\lvert u\rvert}\Gamma[\eta,\underline{\eta}]\Gamma[\mbox{tr}\underline{\chi}]\Gamma[\Psi_{1}]\lesssim\frac{a}{\lvert u\rvert}\lVert{(a^{\frac{1}{2}}\nabla)^{i}(\beta^{R},\underline{\beta}^{R})}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\cdot 1\cdot(\mathcal{R}[\alpha]+1)\\ \lesssim&(\mathcal{R}[\alpha]+1)^{2}\lesssim 1,\end{split}$
@@
where we have used the improvements on $\hat{\chi}$ from Proposition 5.12 and the energy estimate from Proposition 6.6. The rest of the terms can also be bounded above by $1$, using the same approach. We finally arrive at an estimate of the form
@@
From here, collecting all the terms together and using Proposition 6.5, we arrive at the desired result. ∎
This completes the central part of this work. In particular, we prove that there exists an open set of initial data that is almost compatible with the scaling of [3] (with some differences handled by new scaling) and is capable of producing a semi-global characteristic development of the vacuum Einstein’s equations. Simultaneously, we also obtain concentrated generalized Yau mean curvature $c$ (as defined in Theorem 1.2) along the null hypersurface u=au=-a. The idea then is to show that the radius of the boundary $S_{-a,\epsilon}$ is large enough (compared to that of $S_{-a,0}$) so that the condition stated in the theorem 1.2 is met. The last part involves constructing the data on the Cauchy slice $\mathcal{M}_{t=-a}$ and evolving it for a short enough time $\epsilon>0$ (see the diagram 1 for clarity). The motivation behind solving the characteristic initial value problem is to concentrate the mean curvature along the incoming direction by means of large conjugate shear $\underline{\hat{\chi}}$ concentrated along the initial null-hypersurface $\underline{H}_{0}$ starting from dispersed data at past null-infinity.
7. The Cauchy problem on the slab $a,-a+\epsilon]\times\widehat{\mathcal{M}}$ and completion of the argument
In this section, we complete the proof of Theorem 3.1 by establishing the formation of an apparent horizon strictly to the future of the Cauchy hypersurface
$\widehat{\mathcal{M}}_{t=-a}\subset\widehat{\mathcal{M}}$.
We work in spacetime harmonic gauge and solve the vacuum Einstein equations on the slab
$a,-a+\epsilon]\times\widehat{\mathcal{M}},\qquad 0<\epsilon\ll 1,$
with initial data assembled by a gluing procedure. We then propagate the quasi-local boundary geometry forward in time, ultimately showing that the Schoen–Yau curvature radius condition is violated in the future domain, forcing the existence of an apparent horizon. In this section, we provide a sketch of the construction of the data. As mentioned previously, the detailed construction is executed in the sequel [28] since it is, by itself, a very nontrivial problem to solve.
@@ 7.1. Geometric setup and decomposition of the initial slice
Let (^,g^)($\widehat{\mathcal{M}}$$\widehat${g}) be the Lorentzian manifold constructed in previous sections. Let $u,\underline{u}$ be double-null coordinates associated to the canonical foliation, and let
@@
denote the future-directed unit timelike vector field orthogonal (in the induced sense) to the leaves $S_{u,\underline{u}}\subset\widehat{\mathcal{M}}_{t=u+\underline{u}}$.
The interior region $\widehat{\mathcal{M}}_{\mathrm{int}}\subset\widehat{\mathcal{M}}$ admits a foliation by topological 22–spheres outside a compact subset. Our analysis requires only the boundary geometry $\partial\widehat{\mathcal{M}}_{\mathrm{int}}$ and its associated Yau radius.
@@
$t_{1}=-a,\qquad t_{2}=-a+\epsilon,$
where $\epsilon>0$ is sufficiently small. Although we do not require the double-null foliation to describe the interior evolution, it is convenient near $\partial\widehat{\mathcal{M}}_{\mathrm{int}}$.
On the initial slice $\widehat{\mathcal{M}}_{t=-a}$ we introduce the decomposition
(194) $\widehat{\mathcal{M}}_{t=-a}=\mathcal{M}_{1}\cup\mathcal{M}_{2}\cup\mathcal{M}_{3},$
@@
$\mathcal{M}_{1}:=\widehat{\mathcal{M}}_{\mathrm{int}},\qquad\mathcal{M}_{2}:=\widehat{\mathcal{M}}_{t=-a}\cap\mathcal{D}_{-a,\epsilon},\qquad\mathcal{M}_{3}:=\widehat{\mathcal{M}}_{t=-a}\setminus(\mathcal{M}_{1}\cup\mathcal{M}_{2}).$

### v4

@@ 5.1. Energy Estimates for the Weyl/Riemann Curvature components
(212) $\begin{split}&a^{\frac{1}{2}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\alpha}\rVert_{L^{2}_{(sc)}(H_{u}^{(0,\underline{u})})}+a^{\frac{1}{2}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\beta}\rVert_{L^{2}_{(sc)}(\underline{H}_{\underline{u}}^{(u_{\infty},u)})}\\ \lesssim&a^{\frac{1}{2}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\alpha}\rVert_{L^{2}_{(sc)}(H_{u_{\infty}}^{(0,\underline{u})})}+a^{\frac{1}{2}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\beta}\rVert_{L^{2}_{(sc)}(\underline{H}_{0}^{(u_{\infty},u)})}\\ &+a^{\frac{1}{2}}\int_{0}^{\underline{u}}\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert}\lVert{a^{\frac{i}{2}}P_{i}\cdot(a^{\frac{1}{2}}\nabla)^{i}\alpha}\rVert_{L^{1}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}u^{\prime}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\\ &+a^{\frac{1}{2}}\int_{0}^{\underline{u}}\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert}\lVert{a^{\frac{i}{2}}Q_{i}\cdot(a^{\frac{1}{2}}\nabla)^{i}\beta}\rVert_{L^{1}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}u^{\prime}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\\ \lesssim&a^{\frac{1}{2}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\alpha}\rVert_{L^{2}_{(sc)}(H_{u_{\infty}}^{(0,\underline{u})})}+a^{\frac{1}{2}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\beta}\rVert_{L^{2}_{(sc)}(\underline{H}_{0}^{(u_{\infty},u)})}+\frac{1}{a^{\frac{1}{2}}}.\end{split}$
@@
Let $(\mathcal{M},g)$ be a vacuum spacetime equipped with a regular double–null foliation $(u,\underline{u})$ and associated null frame $(e_{3},e_{4},e_{A})$, and assume the bootstrap assumptions (118) and the curvature and Ricci coefficient bounds stated in Theorem 2.2. Let $N\geq 0$ be the top derivative order in the energy hierarchy.
Consider any adjacent pair of Weyl curvature components
$(\Psi_{1},\Psi_{2})\in\Big\{a^{\frac{1}{2}}(\beta,(\rho,\sigma)),\;a^{\frac{1}{2}}((\rho,\sigma),\underline{\beta}),\;a^{\frac{1}{2}}(\underline{\beta},\underline{\alpha})\Big\},$
where the factor $a^{1/2}$ denotes the renormalized scaling used in the curvature energy norms. Then for every integer $0\leq i\leq N+4$ there holds the flux bound
(213) $\lVert{\nabla^{i}\Psi_{1}}\rVert_{L^{2}_{(sc)}(H_{u}^{(0,\underline{u})})}+\lVert{\nabla^{i}\Psi_{2}}\rVert_{L^{2}_{(sc)}(\underline{H}_{\underline{u}}^{(u_{\infty},u)})}\;\lesssim\;\lVert{\nabla^{i}\Psi_{1}}\rVert_{L^{2}_{(sc)}(H_{u_{\infty}}^{(0,\underline{u})})}+\lVert{\nabla^{i}\Psi_{2}}\rVert_{L^{2}_{(sc)}(\underline{H}_{0}^{(u_{\infty},u)})}+1.$
The implicit constant depends only on $N$ and is independent of $a$, $u$, and $\underline{u}$.
@@
$\nabla_{3}\Psi_{1}+\left(\frac{1}{2}+s_{2}(\Psi_{1})\right)\mbox{tr}\underline{\chi}\Psi_{1}-D\Psi_{2}=$ $(\psi,\hat{\chi})\Psi,$
(214) $\begin{split}\nabla_{4}\Psi_{2}-\prescript{*}{}{D}\Psi_{1}=(\psi,\underline{\hat{\chi}})(\Psi_{u},\alpha).\end{split}$
@@
(215) $\begin{split}&\nabla_{3}\nabla^{i}\Psi_{1}+\left(\frac{i+1}{2}+s_{2}(\Psi_{1})\right)\mbox{tr}\underline{\chi}\nabla^{i}\Psi_{1}-D\nabla^{i}\Psi_{2}\\ =&\sum_{i_{1}+i_{2}+i_{3}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}+1}\Psi_{2}+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\psi_{g},\hat{\chi})\nabla^{i_{4}}\Psi\\ &+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\underline{\hat{\chi}},\widetilde{\mbox{tr}\underline{\chi}})\nabla^{i_{4}}\Psi_{1}+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}\mbox{tr}\underline{\chi}\nabla^{i_{4}}\Psi_{1}\\ &+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-2}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}(\underline{\hat{\chi}},\mbox{tr}\underline{\chi})\nabla^{i_{4}}\Psi_{1}:=P_{i}.\end{split}$
@@
(216) $\begin{split}&\nabla_{4}\nabla^{i}\Psi_{2}-\prescript{*}{}{D}\nabla^{i}\Psi_{1}\\ =&\sum_{i_{1}+i_{2}+i_{3}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}+1}\Psi_{1}+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\psi_{g},\underline{\hat{\chi}})\nabla^{i_{4}}(\Psi,\alpha)\\ &+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}\Psi_{2}\\ &+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-2}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}\Psi_{2}:=Q_{i}.\end{split}$
@@
(224) $\begin{split}&\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\left(\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-2}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}(\underline{\hat{\chi}},\mbox{tr}\underline{\chi})\nabla^{i_{4}}\Psi_{1}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{\frac{1}{2}}\hskip 1.42262pt\text{d}u^{\prime}\\ \lesssim&\frac{a}{\lvert u\rvert}\Gamma[\mbox{tr}\underline{\chi}]\Gamma[\alpha]^{2}\lesssim\frac{a}{\lvert u\rvert}\Gamma[\eta,\underline{\eta}]\Gamma[\mbox{tr}\underline{\chi}]\Gamma[\Psi_{1}]\lesssim\frac{a}{\lvert u\rvert}\lVert{(a^{\frac{1}{2}}\nabla)^{i}(\beta^{R},\underline{\beta}^{R})}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\cdot 1\cdot(a^{-\frac{1}{2}}\mathcal{R}[\alpha]+1)\\ \lesssim&(a^{-\frac{1}{2}}\mathcal{R}[\alpha]+1)^{2}\lesssim 1,\end{split}$
@@
where we have used the improvements on $\hat{\chi}$ from Proposition 4.11 and the energy estimate from the Proposition 5.6. The rest of the terms can also be bounded above by $1$, using the same approach. We finally arrive at an estimate of the form
@@
From here, collecting all the terms we arrive at the desired result. This concludes the characteristic development. ∎
6. Cauchy Development and Completion of the argument
In this section, the remaining two major tasks are executed. First, we prove a local Cauchy development of the initial data on $\mathcal{M}_{-a}:=\mathcal{M}_{1}\cup\mathcal{M}_{2}\cup\mathcal{M}_{ext}$. Most importantly, the task is prove that the interior estimates on $\mathcal{M}_{1}$ can be propagated up to $O(1)$ time without substantial distortion. In particular, the interior thickness should not be reduced too much while measured in terms of the $H$-radius. This is due to the choice of initial data in the interior. The remarkable point to note here is that this interior data is chosen to be consistent with the induced Cauchy data on $\mathcal{M}_{2}$ by the characteristic development $D_{a,1}$. This choice causes the minimal distortion of the induced Riemannian geometry on $\mathcal{M}_{1}$ under evolution for $O(1)$ time. In effect, the choice of initial data for the Characteristic development is vital in the full analysis. Nevertheless, once one understands the Characteristic development completely, one can in a sense forget about the characteristic problem and start prescribing data on the Cauchy slice $\mathcal{M}_{1}\cup\mathcal{M}_{2}\cup\mathcal{M}_{ext}$ altogether with appropriate asymptotically flat solution for the $\mathcal{M}_{ext}$ end.
@@ 6.1. Geometric setup and decomposition of the initial slice
Let (,g^)($\mathcal{M}$$\widehat${g}) be the Lorentzian manifold constructed in previous sections. Let $u,\underline{u}$ be double-null coordinates associated to the canonical foliation, and let
@@
denote the future-directed unit timelike vector field orthogonal (in the induced sense) to the leaves $S_{u,\underline{u}}\subset\mathcal{M}_{t=u+\underline{u}}$.
The interior region $\mathcal{M}_{1}\subset\mathcal{M}$ admits a foliation by topological 22–spheres outside a compact subset. Our analysis requires only the boundary geometry $\partial\mathcal{M}_{1}$ and its associated Yau radius.
@@
$t_{1}=-a,\qquad t_{2}=-a-1/a+\epsilon,$
where $\epsilon$ will be chosen to be $\frac{3}{4}$. Although we do not require the double-null foliation to describe the interior evolution (and not well defined for the interior), it is convenient near $\partial\mathcal{M}_{1}$.
On the initial slice $\mathcal{M}_{t=-a}$ we introduce the decomposition
(229) $\mathcal{M}_{t=-a}=\mathcal{M}_{1}\cup\mathcal{M}_{2}\cup\mathcal{M}_{3},$
@@
$\mathcal{M}_{1}:=\mathcal{M}_{1},\qquad\mathcal{M}_{2}:=\mathcal{M}_{t=-a}\cap D_{a,1},\qquad\mathcal{M}_{3}:=\mathcal{M}_{t=-a}\setminus(\mathcal{M}_{1}\cup\mathcal{M}_{2})=\mathcal{M}_{ext}.$